A particle is thrown vertically upward. Its velocity at half of the height is 10m/s. Find the maximum height attained by it if g=10m/s2.

Grade:11

3 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear abhingya

let intitial velocity is U and maximum hight is h

so at maximum hight velocity will be zero

0 =U2 -2gh

U2 =2gh ..........1

now given velocity is = 10m/sec at hight h/2

so 102 = U2 -2gh/2

     100 =2gh -gh

    h = 100/10

   h =10 m/sec

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very

quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

shubhamrana07676@gmail.com
5 years ago
The above and is correct but the and has a silly mistake in the last h =10 m/ s it should be h=10 m and was this a real iitian plz ans
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
Let intitial velocity is U and maximum hight is h
so at maximum hight velocity will be zero
0 = U2 – 2gh
U2 = 2gh ..........1
now given velocity is = 10m/sec at hight h/2
so 102 = U2 – 2gh/2
     100 = 2gh – gh
      h = 100/10
      h = 10 m
 
Thanks and regards,
Kushagra